OpenStudy (anonymous):

The probability that an elevator in marcias building will get stuck is 0.035 .if she rides the elevator 1,040 times a year ,how many times can she expect to get stuck in the elevator ? Round your answer to the nearest number of times . a. 364 b.365 c.35 d.36 ?

OpenStudy (anonymous):

for this one should i subtract or what ? not sure here /

OpenStudy (anonymous):

36 multiply

OpenStudy (anonymous):

You multiply .035(1040) and its about 36 times. Pretty crappy elevator.

hero (hero):

is/of = %/100 x/1040 = 3.5/100 100x = 3640 x = 3640/100 x = 36.4 or just 36

OpenStudy (anonymous):

right